OpenStudy (anonymous):

Will Fan And Medal. What is the solution to the equation? (x+8)-14= -15 x=___________

OpenStudy (texaschic101):

x + 8 - 14 = -15 x - 6 = -15 x = -15 + 6 x = - 9 check.. x + 8 - 14 = -15 -9 + 8 - 14 = -15 -23 + 8 = -15 -15 = -15 (correct)